You invest $1000 at 7% per year, compounded annually for 5 years.

A) Will this effectively protect the purchase power of the original principal, given an annual inflation rate of 6%?

B) If so by how much?
Answer #1

Investment amount = 1000

i = 7%

t = 5 yrs

a) Real interest rate = Nominal interest rte - Inflation

= 7% - 6% = 1%

As real interest rate is positive, therefore it will protect the purchase power of the original principal

b) Purchasing power of amount invested = 1000 * (1+0.01)^5 = 1000 * 1.01^5 = 1051.01 = 1051

Diff = 1051 - 1000 = 51